# Mingchen Li1, Jiatan Huang1, Jeremy Yeung1, Anne Blaes2 Steven Johnson3, Hongfang Liu4, Hua Xu5, Rui Zhang1

### v2

@@ Abstract
Medical Large Language Models (LLMs) such as ClinicalCamel 70B, Llama3-OpenBioLLM 70B have demonstrated impressive performance on a wide variety of medical NLP task. However, there still lacks a large language model (LLM) specifically designed for cancer domain. Moreover, these LLMs typically have billions of parameters, making them computationally expensive for healthcare systems. Thus, in this study, we propose CancerLLM, a model with 7 billion parameters and a Mistral-style architecture, pre-trained on 2,676,642 clinical notes and 515,524 pathology reports covering 17 cancer types, followed by fine-tuning on two cancer-relevant tasks, including cancer phenotypes extraction, and cancer diagnosis generation. Our evaluation demonstrated that CancerLLM achieves state-of-the-art results compared to other existing LLMs, with an average F1 score improvement of 7.61%. Additionally, CancerLLM outperforms other models on two proposed robustness testbeds. This illustrates that CancerLLM can be effectively applied to clinical AI systems, enhancing clinical research and healthcare delivery in the field of cancer.
@@ 1 Introduction
Recently, large language models (LLMs) like GPT-4111https://openai.com/index/gpt-4/, and LLama2 (Touvron et al., 2023) have become the dominant technology in various natural language processing (NLP) tasks, these models develop impressive capabilities in different NLP tasks, including information extraction (e.g., (Li & Zhang, 2023; Li et al., 2024a; c; Zhang et al., 2024)), link prediction (e.g., (Li et al., 2024b; Liu et al., 2024; Li et al., 2022)), and question answering (e.g., (Zhuang et al., 2024; Jiang et al., 2024)). While much attention has been given to these models’ capabilities in the general domain, it’s evident that specialist models have the potential to significantly help clinical and biomedical research Singhal et al. (2023). To tailor LLMs for the clinical domain, several specialized models have been developed, such as ClinicalCamel 70B (Toma et al., 2023) and Llama3-OpenBioLLM 70B [^2] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B. Improving domain-specific language models will accelerate breakthroughs in clinical and biomedical research, leading to enhanced patient care.
@@
Despite showcasing impressive general capabilities, current medical LLMs face significant challenges when applied to cancer. This is primarily due to a deficiency in cancer-specific knowledge within these models. This lack of specialized understanding poses a barrier to the effective utilization of LLMs in assisting doctors with cancer diagnosis and treatment planning. Expanding cancer-specific knowledge within LLMs is imperative for surmounting these obstacles and harnessing their full potential in the cancer field. By integrating comprehensive information encompassing aspects like cancer diagnosis, subjective information, objective information, nursing review of systems (nursing ros) records, laboratory test results, patient self-descriptions, medical history, treatment modalities, assessments, and patient outcomes, LLMs can furnish healthcare professionals with more precise and personalized recommendations. This augmented capability holds the promise of revolutionizing cancer care and aiding medical professionals in enhancing diagnostic accuracy and formulating treatment plans. To the best of our knowledge, CancerBERT (Zhou et al., 2022) is the only model specifically designed for the cancer domain. While CancerBERT (Zhou et al., 2022) stands out as a specialized model for cancer-related tasks, its focus is primarily on breast cancer, leaving other cancer types relatively unaddressed. This limitation underscores the need for further development of specialized models that cover a broader spectrum of cancer types and applications. By expanding the scope of such models to encompass various cancer types and different applications, healthcare professionals can access more comprehensive and tailored resources to enhance patient care in the cancer domain.
@@
Figure 1: The evolution of medical LLM performance on cancer phenotype extraction, and diagnosis generation is measured using the average F1 score, which includes Exact Match, BLEU-2, and ROUGE-L. Our CancerLLM achieves the highest performance with an F1 score of 74.54%.
Bridging these gaps, in this work, we introduce CancerLLM, a state-of-the-art LLM with 7 billion parameters designed to elevate LLM’s proficiency in the cancer domain, and aid medical professionals in cancer phenotype extraction, cancer diagnosis generation. Our model represents an important step towards democratizing advancements in the cancer domain of LLMs. More specifically, our research focuses on four steps: 1) pre-training the LLM in the style of Mistral. We pre-train the Mistral 7B using 2,676,642 cancer clinical notes and 515,524 pathology reports from the University of Minnesota (UMN) Clinical Data Repository. 2) fine-tuning the LLM on a specialized dataset. We proposed three datasets focusing on phenotype extraction, and diagnosis generation. Following this, we fine-tuned CancerLLM using instruction learning for the aforementioned tasks. 3) We evaluate the generation ability of CancerLLM against other medical LLMs using Exact Match (Rajpurkar et al., 2016), BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016) metrics. 4) To further evaluate the model’s ability, we propose two robustness testbeds which include counterfactual robustness and misspellings robustness in the special tasks. As shown in Figure  1, our model not only outperforms existing 7B and 13B models by a significant margin but also delivers good results with those of 70B models.
@@
  • We proposed CancerLLM, a large language model which focuses on the cancer domain.
@@
  • We have developed three datasets aimed at fine-tuning CancerLLM for its two downstream tasks.
  • A large in-depth quantitative analysis of the model’s generation ability and robustness on proposed two testbeds.
2 CANCERLLM
Figure 2: Overview of CancerLLM
Figure 2 illustrates the workflow of the CancerLLM, beginning with the injection of cancer knowledge and followed by supervised instruction tuning. Subsequently, we assess the generative capability of our framework compared to current medical LLMs using three evaluation metrics. Additionally, we introduce two testbeds to evaluate the robustness of the LLMs. Finally, the trained CancerLLM is applied to three specific tasks.
2.1 Pre-training Data for Cancer
2.2 Instruction Tuning Data for Cancer
To adapt CancerLLM for various downstream applications, we employ instruction tuning for three distinct tasks: cancer phenotypes extraction, cancer diagnosis generation, and cancer treatment plan generation. This section provides detailed information on each task description and its instruction tuning data.
2.2.1 Cancer Phenotypes extraction
question
1 What is the tumor size in the given context?
2 What is the histological type in the given context?
3 Please identify the receptors mentioned in the provided context.
4 What is the receptor type in the given context?
6 What is the stage of cancer in the given context?
7 Please describe the tumor location in the given context
8 What is the grade of cancer in the given context?
Table 1: Question types
2.2.2 Diagnosis Generation
2.3 Training Methodology
2.3.1 Continued Pre-training
2.3.2 Instruction Tuning
After continued pre-training, our model undergoes instruction tuning with three datasets, aligning with different clinical requirements. We also use LoRA to fine-tune the model. The rank is set to 64, and alpha is set to 16. The training step is 5. We use the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 2e-4. The model was fine-tuned using a single A100 GPU with 4 hours of training time and 7338M of GPU memory.
2.4 Baselines and Evaluation Metrics
In our work, we used 14 widely used clinical, biomedical, and general LLMs as the baseline models, including: six 7B LLMs ( PMC LLaMA 7B Wu et al. (2023), Medalpaca 7B Han et al. (2023), LLAMA-2 7B Han et al. (2023), Mistral 1*7B Jiang et al. (2023), Mixtral 8*7B Jiang et al. (2023) , and Bio-Mistral 7B Labrak et al. (2024) ), one 8B LLM ( LLama3 8B [^4] https://huggingface.co/meta-llama/Meta-Llama-3-8B), four 13B models (MedLLaMA 13B Wu et al. (2023) , PMC LLaMA 13B Wu et al. (2023), Medalpaca 13B Han et al. (2023), and LLaMA2 13B Touvron et al. (2023) ) three 70B LLMs ( LLaMA2 70B Touvron et al. (2023) , Llama3-OpenBioLLM-70B [^5] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B and ClinicalCamel-70B Toma et al. (2023) ). Same with  Li & Huang (2023), we evaluate all the models based on generative evaluation metrics, including Exact Match (Rajpurkar et al., 2016) , BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016).
2.5 Robustness Testbeds
Counterfactual Robustness
[…]
Misspellings Robustness
[…]
@@ 3.1.1 Cancer Diagnosis Generation
To assess the scalability of our model, in this part, we evaluate the performance of CancerLLM in diagnosis generation task. Table 2 presents the experiment results of various medical LLMs. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 28.93%, and 17.92% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (3) Our model not only achieves the best performance among 7B models but also surpasses the performance of medical LLMs with 13B and 70B parameters. (4) Without domain knowledge, larger parameter LLMs are ineffective. For instance, Llama3-OpenBioLLM-70B and ClinicalCamel-70B achieve only 62.81% and 63.55%, separately. (5) Table 3 presents the experiment results of various medical LLMs on ICDdiagnosis, We observed that CancerLLM improve the best best baseline model Bio-Mistral 7B by 4.36% in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (6) The 70B models do not achieve the best performance.
Exact Match BLEU-2 ROUGE-L F1
PMC LLaMA 7B 24.33 24.33 24.33 35.66 35.66 35.66 44.66 44.66 44.66 34.90
Medalpaca 7B 27.27 27.27 27.27 38.09 38.09 38.09 47.86 47.86 47.86 37.73
LLAMA-2 7B 24.06 24.06 24.06 35.02 35.02 35.02 44.36 44.36 44.36 34.48
Mistral 1*7B 26.74 26.74 26.74 39.00 39.00 39.00 48.53 48.53 48.53 38.09
Mixtral 8*7B 29.68 29.68 29.68 38.18 38.18 38.18 47.86 47.86 47.86 38.57
Bio-Mistral 7B 30.21 30.21 30.21 41.16 41.16 41.16 50.69 50.69 50.69 40.68
LLama3 8B 27.01 27.01 27.01 38.16 38.16 38.16 48.05 48.05 48.05 37.73
MedLLaMA 13B 29.41 29.41 29.41 39.80 39.80 39.80 48.34 48.34 48.34 39.18
PMC LLaMA 13B 26.74 26.74 26.74 37.02 37.02 37.02 46.03 46.03 46.03 36.59
Medalpaca 13B 25.13 25.13 25.13 35.57 35.57 35.57 45.31 45.31 45.31 35.33
LLaMA2 13B 22.99 22.99 22.99 34.08 34.08 34.08 43.42 43.42 43.42 33.49
LLaMA2 70B 27.81 27.81 27.81 38.85 38.85 38.85 47.94 47.94 47.94 38.20
Llama3-OpenBioLLM-70B 27.01 27.01 27.01 39.39 39.39 39.39 49.35 49.35 49.35 38.58
ClinicalCamel-70B 28.34 28.34 28.34 38.75 38.75 38.75 48.05 48.05 48.05 38.37
CancerLLM(Ours) 34.49 34.49 34.49 46.03 46.03 46.03 54.62 54.62 54.62 45.04
Table 3: Results of Cancer Diagnosis Generation on ICDdiagnosis
@@ 3.2.2 Misspellings Robustness
ICDdiagnosis Bio-Mistral 7B 2% 0.00 0.00 0.00 7.07 7.07 7.07 18.00 18.00 18.00 8.62
4% 0.00 0.00 0.00 7.27 7.27 7.27 17.78 17.78 17.78
6% 0.27 0.27 0.27 8.07 8.07 8.07 18.63 18.63 18.63
8% 0.00 0.00 0.00 7.65 7.65 7.65 18.79 18.79 18.79
CancerLLM(Ours) 2% 0 0 0 8.93 8.93 8.93 20.59 20.59 20.59 9.57
4% 0.27 0.27 0.27 9.07 9.07 9.07 20.53 20.53 20.53
6% 0.27 0.27 0.27 8.34 8.34 8.34 19.34 19.34 19.34
8% 0 0 0 8.20 8.20 8.20 19.31 19.31 19.31
Table 6: Misspellings robustness performance on diagnosis generation and ICDdiagnosis. The rate refers to different misspelling rates in each instance.
To validate the robustness of our model, we propose a misspelling robustness testbed designed to simulate misspelling in the clinical notes. Table 6 presents the complete results and average F1 performance of different counterfactual rates across exact match, BLEU-2, and ROUGE-L metrics for both diagnosis generation and ICDdiagnosis. Our observations are as follows: CancerLLM consistently outperforms Bio-Mistral 7B. For the diagnosis generation task, CancerLLM achieves a similar average F1 score compared to Bio-Mistral 7B. Specifically, in ICD diagnosis, CancerLLM scores 9.57%, while Bio-Mistral 7B scores 8.62%.
@@ 3.3 Results of Generation Time and GPU Memory
Phenotype Extraction Diagnosis Generation1
@@ 4.1 Cancer Diagnosis Generation
In this task, we primarily pre-train CancerLLM and explore its effectiveness in the diagnosis generation task by providing relevant information. As shown in Table 2, the Bio-Mistrial 7B obtains the best performance among all baseline LLMs. We guess that one reason for this is that Bio-Mistral 7B is trained on a large pre-processed corpus, such as the PubMed Central corpus Canese & Weis (2013). Another reason is that the foundational structure of Mixtral enhances training effectiveness. The Mixtral 8*7B outperforms the Mixtral 1*7B by more than 5% in average F1 value. However, Mixtral 8*7B is a Mixture of Experts (MoE) model with 8 experts, making training and inference time challenging. Fortunately, our CancerLLM achieves the best model performance through pre-training and fine-tuning with a single expert model. Not all LLMs with large parameters achieve the best performance. For example, the 70B models (LLaMA2 70B, Llama3-OpenBioLLM-70B, and ClinicalCamel-70B) do not outperform the Bio-Mistral 7B, despite being trained on medical corpora. This indicates that smaller parameter LLMs can achieve better performance. For ICDdiagnosis, we find that overall performance is lower than in diagnosis generation. This is because the expected diagnoses in ICD diagnosis are more complex. Our 7B model, CancerLLM, which incorporates cancer domain knowledge, also supports this point. This is particularly important in the medical domain as it reduces training time and memory usage, providing opportunities for medical institutions with limited computational resources.

### v3

# Mingchen Li1, Zaifu Zhan2, Jiatan Huang1, Jeremy Yeung1, Kai Ding3, Anne Blaes4 Steven Johnson5, Hongfang Liu6, Hua Xu7, Rui Zhang1

@@ Abstract
Medical Large Language Models (LLMs) have demonstrated impressive performance on a wide variety of medical NLP tasks; however, there still lacks a LLM specifically designed for phenotyping identification and diagnosis in cancer domain. Moreover, these LLMs typically have several billions of parameters, making them computationally expensive for healthcare systems. Thus, in this study, we propose CancerLLM, a model with 7 billion parameters and a Mistral-style architecture, pre-trained on nearly 2.7M clinical notes and over 515K pathology reports covering 17 cancer types, followed by fine-tuning on two cancer-relevant tasks, including cancer phenotypes extraction and cancer diagnosis generation. Our evaluation demonstrated that the CancerLLM achieves state-of-the-art results with F1 score of 91.78% on phenotyping extraction and 86.81% on disganois generation. It outperformed existing LLMs, with an average F1 score improvement of 9.23%. Additionally, the CancerLLM demonstrated its efficiency on time and GPU usage, and robustness comparing with other LLMs. We demonstrated that CancerLLM can potentially provide an effective and robust solution to advance clinical research and practice in cancer domain.
@@ 1 Introduction
Recently, large language models (LLMs) like GPT-4111https://openai.com/index/gpt-4/, and LLama2 (Touvron et al., 2023) have become the dominant technology in various natural language processing (NLP) tasks, these models develop impressive capabilities in different NLP tasks, including information extraction (e.g., (Li & Zhang, 2023; Li et al., 2024a; c; Zhang et al., 2024)), link prediction (e.g., (Li et al., 2024b; Liu et al., 2024; Li et al., 2022)), and question answering (e.g., (Zhuang et al., 2024; Jiang et al., 2024; Huang et al., 2024)). While much attention has been given to these models’ capabilities in the general domain, it’s evident that specialist models have the potential to significantly help clinical and biomedical research Singhal et al. (2023). To tailor LLMs for the clinical domain, several specialized models have been developed, such as ClinicalCamel 70B (Toma et al., 2023) and Llama3-OpenBioLLM 70B [^2] https://huggingface.co/aaditya/Llama3-OpenBioLLM-70B. Improving domain-specific language models will accelerate breakthroughs in clinical and biomedical research, leading to enhanced patient care.
@@
Despite showcasing impressive general capabilities, current medical LLMs face significant challenges when applied to cancer (Yang et al., 2023; Ullah et al., 2024; Zhou et al., 2024). This is primarily due to a deficiency in cancer-specific knowledge within these models (Perez-Lopez et al., 2024; Truhn et al., 2023). This lack of specialized understanding poses a barrier to the effective utilization of LLMs in assisting doctors with cancer diagnosis and treatment planning. Expanding cancer-specific knowledge within LLMs is imperative for surmounting these obstacles and harnessing their full potential in the cancer field (Tsang et al., 2025). By integrating comprehensive information encompassing aspects like cancer diagnosis, subjective information, objective information, nursing review of systems (nursing ros) records, laboratory test results, patient self-descriptions, medical history, treatment modalities, assessments, and patient outcomes, LLMs can furnish healthcare professionals with more precise and personalized recommendations (AlSaad et al., 2024). This augmented capability holds the promise of revolutionizing cancer care and aiding medical professionals in enhancing diagnostic accuracy and formulating treatment plans. To the best of our knowledge, CancerBERT (Zhou et al., 2022) is the only model specifically designed for the cancer domain. While CancerBERT (Zhou et al., 2022) stands out as a specialized model for cancer-related tasks, its focus is primarily on breast cancer, leaving other cancer types relatively unaddressed. This limitation underscores the need for further development of specialized models that cover a broader spectrum of cancer types and applications. By expanding the scope of such models to encompass various cancer types and different applications, healthcare professionals can access more comprehensive and tailored resources to enhance patient care in the cancer domain.
Except for the lack of knowledge, to evaluate the capability of LLMs in the cancer domain, the lack of high-quality datasets grounded in real-world electronic health records (EHRs) remains a major challenge (Zaied et al., 2015; Dar et al., 2022). Most existing EHR datasets are proprietary and restricted due to privacy concerns, limiting the community’s ability to systematically evaluate LLMs on clinically meaningful tasks. As such, constructing dedicated datasets tailored to cancer-specific applications is essential for enabling accurate, task-relevant assessment of model capabilities (Varlamova et al., 2024; Vorontsov et al., 2024). These datasets play a pivotal role in benchmarking LLM performance under realistic clinical conditions and guiding the development of domain-adapted models that can truly support oncology care.
@@
Figure 1: The evolution of medical LLM performance on cancer phenotype extraction, and diagnosis generation is measured using the average F1 score, which includes Exact Match, BLEU-2, and ROUGE-L. Our CancerLLM achieves the highest performance with an F1 score of 89.30%.
To bridge these gaps, we first construct two datasets focused on cancer phenotype extraction and diagnosis generation. We then introduce CancerLLM, a state-of-the-art 7B-parameter language model tailored to enhance LLM proficiency in the cancer domain and support medical professionals in both phenotype extraction and diagnosis generation tasks. Our model represents an important step towards democratizing advancements in the cancer domain of LLMs. More specifically, our research focuses on four steps: 1) pre-training the LLM in the style of Mistral. We pre-train the Mistral 7B using 2,676,642 cancer clinical notes and 515,524 pathology reports from the University of Minnesota (UMN) Clinical Data Repository. 2) fine-tuning the LLM on a specialized dataset. We proposed two datasets focusing on phenotype extraction and diagnosis generation to evaluate the information extraction and diagnosis ability of CancerLLM. Following this, we fine-tuned CancerLLM using instruction learning for the aforementioned tasks. 3) We evaluate the generation ability of CancerLLM against other medical LLMs using Exact Match (Rajpurkar et al., 2016), BLEU-2 (Papineni et al., 2002), and ROUGE-L (Cohan & Goharian, 2016) metrics. 4) To further evaluate the model’s robustness, we propose two robustness testbeds, which include counterfactual robustness and misspellings robustness in the special tasks. 5) To examine the effectiveness of different examples in instruction tuning, we proposed retrieval-based CanerLLM, and further evaluated the task performance of five retrieval-based models built upon the Cancer LLM.
As shown in Figure  1, our model not only outperforms existing 7B and 13B models by a significant margin but also delivers good results with those of 70B models.
@@
  • We proposed CancerLLM, a large language model which focuses on the cancer domain, as well as retrieval-based CancerLLM with five retrieval-based models.
@@
  • We conducted an in-depth quantitative analysis of the model’s generation ability and robustness on proposed two testbeds on phenotype extraction and diagnosis generation.
@@ 2.1.1 Cancer Diagnosis Generation
To assess the scalability of our model, in this part, we evaluate the performance of CancerLLM in diagnosis generation task. Table 1 presents the experiment results of various medical LLMs. We have the following observations: (1) our CancerLLM significantly outperforms all the strong baselines and its variants across all evaluation metrics. (2) We observed that CancerLLM improve the original Mistral 1*7B, and Bio-Mistral 7B by 31.7%, and 17.92% respectively, in term of average F1 across Exact Match, BLEU–2 and ROUGE-L. (3) Our model not only achieves the best performance among 7B models, but also outperforms medical LLMs with 13B and 70B parameters, demonstrating that a lightweight, domain-specific model can effectively rival much larger general-purpose models.
@@ 2.2.2 Misspellings Robustness
Table 4: Misspellings robustness performance on diagnosis generation. The rate refers to different misspelling rates in each instance.
Table 4 presents the robustness evaluation of CancerLLM and Bio-Mistral 7B under varying levels of misspellings in the diagnosis generation task. The misspelling rates range from 2% to 8%, and the table reports performance using Exact Match, BLEU-2, and ROUGE-L metrics. Both models show a drop in performance as the misspelling rate increases, which is expected. Bio-Mistral 7B achieves slightly higher average F1 scores at lower error rates, especially at 2% and 4%. However, CancerLLM demonstrates more consistent performance across all levels, notably outperforming Bio-Mistral at the 8% rate in terms of Exact Match (0.27% vs. 0.09%) and ROUGE-L (25.81% vs. 23.80%). The overall average F1 scores of both models are very close, with Bio-Mistral 7B at 12.06% and CancerLLM at 12.00%, indicating comparable robustness. These results suggest that CancerLLM maintains stable generation quality under noisy clinical inputs, validating its reliability in real-world settings
@@ 2.3 Results of Generation Time and GPU Memory
Phenotype Extraction Diagnosis Generation
@@
2.4 Results of Retrieval-based CancerLLM
Exact Match BLEU-2 ROUGE-L F1
Task Retriever Precision Recall F1 Precision Recall F1 Precision Recall F1 Average
Phenotype Extraction Random 89.47 89.47 89.47 91.30 91.30 91.30 93.91 93.91 93.91 91.55
Medcpt 87.95 87.95 87.95 90.18 90.18 90.18 92.97 92.97 92.97 90.36
Contriever 88.14 88.14 88.14 90.62 90.62 90.62 94.04 94.04 94.04 90.93
SGPT 83.02 83.02 83.02 85.67 85.67 85.67 89.93 89.93 89.93 86.20
Specter2 88.61 88.61 88.61 90.81 90.81 90.81 93.67 93.67 93.67 91.03
No-retriever 89.37 89.37 89.37 91.98 91.98 91.98 93.98 93.98 93.98 91.78
Diagnosis Generation Random 43.30 43.30 43.30 52.90 52.90 52.90 63.63 63.63 63.63 53.27
Medcpt 58.34 58.34 58.34 65.99 65.99 65.99 73.54 73.54 73.54 65.95
Contriever 82.50 82.50 82.50 85.15 85.15 85.15 88.80 88.80 88.80 85.48
SGPT 43.94 43.94 43.94 52.97 52.97 52.97 63.75 63.75 63.75 53.55
Specter2 85.78 85.78 85.78 89.09 89.09 89.09 92.49 92.49 92.49 89.12
No-retriever 83.50 83.50 83.50 86.60 86.60 86.60 90.34 90.34 90.34 86.81
Table 6: Results of retrieval-based CancerLLM
As shown in Table 6. We evaluated the performance of CancerLLM with different retrievers on two tasks: phenotype extraction and diagnosis generation. For phenotype extraction, all retrievers achieved strong results, with Random and Specter2 performing slightly better than others, but the no-retriever baseline surprisingly achieved the highest average F1 score of 91.78%. Contriever also showed competitive performance with an average F1 of 90.93%, slightly outperforming MedCPT and SGPT. In contrast, diagnosis generation revealed a larger performance gap between retrievers. Specter2 led all models with an average F1 score of 89.12, followed by Contriever and the no-retriever baseline at 85.48% and 86.81%, respectively. Random and SGPT performed the worst in this task, with average F1 scores of only 53.27% and 53.55%. Overall, Specter2 and Contriever consistently delivered strong performance across both tasks, while the no-retriever baseline remained surprisingly competitive, especially in phenotype extraction.
@@ 3.1 Cancer Diagnosis Generation
In this task, we primarily pre-train CancerLLM and explore its effectiveness in the diagnosis generation task by providing relevant information. As shown in Table 1, the Bio-Mistrial 7B obtains the best performance among all baseline LLMs. We guess that one reason for this is that Bio-Mistral 7B is trained on a large pre-processed corpus, such as the PubMed Central corpus Canese & Weis (2013). Another reason is that the foundational structure of Mixtral enhances training effectiveness. The Mixtral 8X7B outperforms the Mixtral 1X7B by more than 5% in average F1 value. However, Mixtral 8X7B is a Mixture of Experts (MoE) model with 8 experts, making training and inference time challenging. Fortunately, our CancerLLM achieves the best model performance through pre-training and fine-tuning with a single expert model. Not all LLMs with large parameters achieve the best performance. For example, the 70B models (LLaMA2 70B, Llama3-OpenBioLLM-70B, and ClinicalCamel-70B) do not outperform the Bio-Mistral 7B, despite being trained on medical corpora. This indicates that smaller parameter LLMs can achieve better performance.